# LLMs Can’t Plan, But Can Help Planning in LLM-Modulo Frameworks

### 02 Feb 2024 v1

@@ 1 Introduction
Large Language Models (LLMs), essentially n-gram models on steroids which have been trained on web-scale language corpora (or, effectively, our collective consciousness), have caught the imagination of the AI research community with linguistic behaviors that no one expected text completion systems to possess. Their seeming versatility has led many researchers to wonder whether they can also do well on planning and reasoning tasks typically associated with System 2 competency. On the face of it, this doesn’t seem to ring true, as both by training and operation, LLMs are best seen as giant pseudo System 1 Kahneman (2011) (see Figure 1). Even from a pure engineering perspective, a system that takes constant time to produce the next token cannot possibly be doing principled reasoning on its own. [^1] Think of asking an LLM an yes/no question–is this theorem logically entailed by this first-order logic knowledge-base. This is well-known to be a semi-decidable problem. Ask yourself if the LLM will take longer in answering the question. (If you are thinking Chain-of-thought prompts or training with step-by-step data, consider that you are essentially changing the nature of the original prompt/training). Not surprisingly, initial excitement based on anecdotal performance of LLMs on reasoning tasks Bubeck et al. (2023) has dissipated to some extent by the recent spate of studies questioning the robustness of such behaviors–be it planning Valmeekam et al. (2023c); Kambhampati (2023), simple arithmetic and logic Dziri et al. (2023), theory of mind abilities Ullman (2023); Verma et al. (2024), or general mathematical and abstract benchmarks McCoy et al. (2023); Gendron et al. (2023). Despite this, a steady stream of claims continue to be made in the literature about the planning and reasoning capabilities of LLMs.
@@
In truth, LLMs are a whole lot more than machine translators. They are a kind of approximate knowledge sources (albeit sans guarantees) trained on our collective consciousness. While it is unlikely that they will have System 2 competencies by themselves, they can nevertheless be valuable resources in solving System 2 tasks. The problem with Alchemy is not that Chemistry is useless, but that people wanted to delude themselves that Chemistry–a pretty amazing discipline on its own merits–can be Nuclear Physics if you prompt it just so. The confusions regarding LLM abilities, or should we say, LLM alchemy, seems to be not that much different–oscillating between ignoring what they are good at, and ascribing them abilities they don’t have.
@@
Second, we will propose a framework that allows us to leverage LLMs effectively in planning tasks, by combining them with external critics, verifiers and humans. We call this an LLM-Modulo Framework (a name loosely inspired by SAT Modulo Theories Nieuwenhuis & Oliveras (2006)); see Figure 3. LLMs play a spectrum of roles in this architecture, from guessing candidate plans, to translating those plans into syntactic forms that are more accessible to external critics, to helping end users flesh out incomplete specifications to helping expert users acquire domain models (that in turn drive model-based critics). All this leveraging of LLMs is done without ascribing to them any planning verification abilities–thus ensuring that the plans generated in this architecture can have formal correctness guarantees where possible.
One question regarding this LLM-Modulo architecture from a planning perspective is whether it is more han a gratuitous attempt to shoe-horn LLMs to solve planning problems, when there are more formal combinatorial planning systems available in multiple communities Ghallab et al. (2004). Compared to a planner that is guaranteed to be correct in a narrow set of domains, LLMs may likely be good at generating plausible (but not guaranteed to be correct) plan heuristics/suggestions in many more scenarios. Thus, unlike the traditional planning architectures studied in AI Ghallab et al. (2004), which put a priori constraints on the expressiveness of the problems that can be posed to the planner (to wit, the different expressiveness levels of the PDDL specification McDermott et al. (1998)), the LLM-Modulo architecture puts no such restrictions. In this sense, it is more representative of real-world planning problems such as those in NASA mission planning, where the different critics–human and automated–are at best able to give “no objection” certificates, clearing it from their perspective. (Indeed, both deep space network planning and mars rover task planning are done via a collective human blackboard. Johnston et al. (2014) Bresina et al. (2004).) Note that this is starkly different from just sending an unvetted plan out to execution (as would be the case if we have LLMs operate in autonomous mode to guess plans). Generalizing planning and reasoning frameworks this way is consistent with the Doyle & Patil’s call to the Knowledge Representation community of yore Doyle & Patil (1991).
@@ 2.1 LLMs cannot generate executable plans in autonomous mode
They show that the results in the autonomous mode are pretty bleak. On average, only about 12% of the plans that the best LLM (GPT-4) generates are actually executable without errors and reach their goals. They show that the choice of the specific LLM (they have tested the family of GPT LLMs including GPT-4 OpenAI (2023), GPT-3.5 OpenAI (2022), InstructGPT-3 Ouyang et al. (2022) and GPT-3 Brown et al. (2020)), as well as fine-tuning does not seem to have a major effect on this dismal performance. They also show that the performance deteriorates further if the names of the actions and objects in the domain are obfuscated–a change that doesn’t in any way affect the performance of the standard AI planners. This latter further suggests that LLMs are more likely doing approximate retrieval of plans than planning.
@@ 2.2 LLMs cannot verify plans and thus cannot improve by self-critiquing
There still exists considerable optimism that even if LLMs can’t generate correct solutions in one go, their accuracy might improve in an iterative prompting regime, where LLMs will be able to ”self-critique” their candidate solutions and refine them to the point of correctness Yao et al. (2023b, a); Shinn et al. (2023); Weng et al. (2023); Huang et al. (2022). This belief seems to rest largely on the assumption that verification of correctness should be easier than generation for many reasoning problems–a rather classical argument from computational complexity. There are grounds to be skeptical of this assumption as the complexity of the reasoning task should be irrelevant to LLM performance if what they are doing is approximate retrieval. In general, unless LLMs are trained not just on “correct data,” but also “corrections data,” there is no a priori reason to believe that their critiques would even be approximately relevant, let alone actually correct.
@@
Their results indicate that in direct mode, LLMs are, perhaps not surprisingly, pretty bad at solving graph coloring instances. More interestingly, they are no better at verifying solutions. In iterative modes, given the inability of LLMs to verify solutions, it should come as no surprise that their experiments also show that the strategy of LLMs self-critiquing their solutions does not improve over the baseline. It is apparently worse because the system can’t recognize a correct coloring and thus merrily passes over fortuitously correct colorings it has generated, ending up with a wrong one! Similar results have also been reported for planning problems in Valmeekam et al. (2023c).
One important corollary of the fact that LLMs cannot self-critique their plans is that they can’t also self-improve by generating synthetic data by generating plans themselves, critiquing the plans by themselves to improve them, and then use those to fine-tune themselves, as has been claimed in the literature Wang et al. (2022) [^3] Contrary to the claim of “self-improvement”, works like Wang et al. (2022) actually heavily depend on external knowledge (crafted seed examples) and critics (filtering step). Huang et al. (2023b); see Section 3.3.
@@ 2.3 Analyzing Claims to the Contrary in the Literature
Claims about Planning: To analyze the planning claims, we need to first understand that solving planning tasks requires (a) having the necessary planning domain knowledge–the actions and their preconditions, effects; the standard hierarchical recipes (e.g. task reduction schemas in HTN planning), past cases/plans, etc., and (b) being able to assemble this planning knowledge into an executable plan that takes care of any subgoal/resource interactions. The first can be called the knowledge acquisition and the second reasoning/planning part. Many of the papers claiming planning abilities of LLMs, on closer examination, wind up confusing general planning knowledge extracted from the LLMs for executable plans. When all we are looking for are abstract plans, such as “wedding plans,” with no intention of actually executing the said plans, it is easy to confuse them for complete executable plans. Indeed, our close examination of several works claiming planning capabilities for LLMs suggests that they either work in domains/tasks where subgoal interactions can be safely ignored Yao et al. (2023b); Shinn et al. (2023) [^4] Although domains like AlfWorld Shridhar et al. (2021) do have sub-goal interactions for successful task completion, Yao et al. (2023b) and Shinn et al. (2023) ignore these interactions relying on the ergodic nature of the domain when prompting LLMs for generating plans. –either because they are just working on a single goal, or because the world is forgiving and ergodic; or delegate the interaction resolution (reasoning) to the humans in the loop (who, through repeated prompting, have to “correct” the plan). Sometimes, in common sense domains, or with enough fine-tuning, the “assembling” part may also be obviated by having seen a case that pretty much corresponds to the problem that needs to be solved. Not surprisingly, the work by Valmeekam et al. (2023c) shows that if the action interactions are removed by relaxing the world models, then the ability of LLMs to guess executable plans improves. Without these assumptions or mitigations, the plans that come out of LLMs may look reasonable to the lay user, and yet lead to execution time interactions and errors. [^5] These issues are illustrated in part by a recent news story Kugel & Hiltner (2023) about the proliferation of travel planning books, mostly auto-extracted from LLMs, and the ensuing disappointment of the unsuspecting end users who buy them mistaking them for usable plans!
@@
The fact that LLMs are often good at extracting planning knowledge can indeed be gainfully leveraged. As shown in recent works Guan et al. (2023), LLMs can be a rich source of approximate models of world/domain dynamics and user preferences, as long as the humans (and any specialized critics) in the loop verify and refine those models, and give them over to model-based solvers. This way of using LLMs has the advantage that the humans need only be present when the dynamics/preference model is being teased out and refined, and the actual planning after that can be left to sounder planning frameworks with correctness guarantees, such as LLM-Modulo.
@@
Such an overall approach has striking similarities to knowledge-based AI systems of yore, with LLMs effectively replacing the “knowledge engineer” (see Figure 2). Given the rather quixotic and dogmatic shift of AI away from approaches that accept domain knowledge from human experts that some writers termed “Polanyi’s Revenge” (c.f. Kambhampati (2021)), this new trend of using LLMs as knowledge sources can be viewed as a form of avenging Polanyi’s revenge! Indeed, LLMs make it easy to get problem-specific knowledge as long as we are willing to relax the correctness requirements of that knowledge. In contrast to the old knowledge engineering approaches, LLMs offer this without making it look like we are inconveniencing any specific human (we are, instead, just leveraging everything humans told each other!). So the million dollar question for reasoning tasks is: ”how would you do robust planning if you have some doddering know-it-all ready to give you any kind of knowledge?” The LLM-Modulo Framework is a principled method for tackling this challenge.
@@
In a related vein, there is the recent Tree of Thoughts (ToT) paper Yao et al. (2023a), which has been pitched as a way to convert LLMs into some type of systematic search with self-verification. A closer look at the work however shows that ToT simply iteratively back-prompts the LLM until it comes up with a solution that is acceptable to an external verifier. Specifically, ToT employs a problem-specific prompt priming method. The “tree” in ToT is essentially a way to generate diverse priming prompts (that the authors set up in a problem specific way). In other words, despite the use of terminology of problem-solving agents Russell & Norvig (2010)–search tree, expansion etc., there is no deeper connection to search-based agents. [^7] Our preliminary experiments also show that at least in 24 puzzle, a simple iterative prompting, even without a systematic prompt diversification, is quite competitive with the ToT framework.
@@
In general planning problems, one way to provide an external verifier is to (a) write a domain model (e.g. in PDDL) and (b) feed it to an off-the-shelf model-based verifier like VAL (c.f. Guan et al. (2023)
@@ 3.1 Critics
Critics can range from those that check for the hard constraints of correctness–causal correctness, timeline correctness, resource constraint correctness etc., to those that check for soft constraints such as style, explicability, preference conformance etc. As discussed in Section 2.3, while LLMs cannot take on the role of hard critics, they can help simulate some soft critics (e.g. Verma et al. (2024)). For PDDL planning problems, the hard critic can be based on VAL Howey et al. (2004), which works off of a model (which itself can be acquired with the help of the LLM Guan et al. (2023).
@@ 3.2 Backprompt (Meta) Controller
The processing steps taken in the controller can range from simple round-robin selection of prompts to generating a summarized prompt (with LLM help) to employ a prompt diversification strategy to elicit the next candidate from a different part of the implicit search space (akin effectively to the strategy used in systems such as Tree of Thoughts prompting Yao et al. (2023a), as discussed in 2.3).
@@ 3.3 Fine Tuning & Synthetic Data
One important question is where this additional data for fine tuning comes from. A tempting idea is to have the LLM itself generate this additional data, and improve it by self-critiquing/verification, before fine-tuning itself on the data. This EM-like approach unfortunately will not work given that LLMs can’t verify their own solutions (see Section 2.2). In the past, this meant that the only reliable way to generate synthetic data is to use external plan generators–for example, use a classical planner like FF to solve a host of blocks world instances and then use those solutions to fine-tune the LLM. The LLM-Modulo framework, in contrast, provides an alternative way of using an LLM-based framework to generate synthetic data that is guaranteed correct.
@@ 3.4 Specification Refinement & Model Acquisition (Semi-automated)
As mentioned earlier, we avoid having humans involved in iteratively prompting LLMs–as this can be an infeasibly time-consuming activity for them. Instead, we let automated verifiers, either model-based or LLM-supported, to manage the plan critiquing process. The framework does depend on humans for “once per domain” and “once per problem” critiques. In the former category, human domain experts can play a role in acquiring the domain model with the help of the LLM. Examples of such interaction include teasing out PDDL planning models from the LLMs with the help of human expert curation (top left in Figure 3). The idea here is that the traditional domain model acquisition task (e.g. sim (2001)) is significantly made easier by having the LLMs help with ideas regarding various pieces of the domain model (e.g., actions, their preconditions and effects) and letting humans sign off/critique the resulting model. Once the model is acquired this way, it can be used by correctness verifiers such as VAL Howey et al. (2004); Guan et al. (2023). Often the planning problems in real world situations are specified incompletely, leaving it to the human commonsense to refine the specification. This brings up a second role for humans–this time end users (bottom left in Figure 3–in collaboratively refining the specification with the help of LLMs (similar to the way done in Xie et al. (2023); Liu et al. (2023)).

### 06 Feb 2024 v2

@@ 1 Introduction
Large Language Models (LLMs), essentially n-gram models on steroids which have been trained on web-scale language corpora (or, effectively, our collective consciousness), have caught the imagination of the AI research community with linguistic behaviors that no one expected text completion systems to possess. Their seeming versatility has led many researchers to wonder whether they can also do well on planning and reasoning tasks typically associated with System 2 competency. On the face of it, this doesn’t seem to ring true, as both by training and operation, LLMs are best seen as a giant pseudo System 1 Kahneman (2011) (see Figure 1). Even from a pure engineering perspective, a system that takes constant time to produce the next token cannot possibly be doing principled reasoning on its own. [^1] Think of asking an LLM an yes/no question–is this theorem logically entailed by this first-order logic knowledge-base. This is well-known to be a semi-decidable problem. Ask yourself if the LLM will take longer in answering the question. (If you are thinking Chain-of-thought prompts or training with step-by-step data, consider that you are essentially changing the nature of the original prompt/training). Not surprisingly, initial excitement based on anecdotal performance of LLMs on reasoning tasks Bubeck et al. (2023) has dissipated to some extent by the recent spate of studies questioning the robustness of such behaviors–be it planning Valmeekam et al. (2023c); Kambhampati (2023), simple arithmetic and logic Dziri et al. (2023), theory of mind abilities Ullman (2023); Verma et al. (2024), or general mathematical and abstract benchmarks McCoy et al. (2023); Gendron et al. (2023). Despite this, a steady stream of claims continue to be made in the literature about the planning and reasoning capabilities of LLMs.
@@
In truth, LLMs are a whole lot more than machine translators. They are a kind of approximate knowledge sources (albeit sans guarantees) trained on our collective consciousness. While it is unlikely that they will have System 2 competencies by themselves, they can nevertheless be valuable resources in solving System 2 tasks. To put it another way, the problem with Alchemy of yore is not that Chemistry is useless, but that people wanted to delude themselves that Chemistry–a pretty amazing discipline on its own merits–can be Nuclear Physics if you prompt it just so. The confusions regarding LLM abilities, or should we say, LLM alchemy, seems to be not that much different–oscillating between ignoring what they are good at, and ascribing them abilities they don’t have.
@@
Second, we will propose a framework that allows us to leverage LLMs effectively in planning tasks, by combining them with external critics, verifiers and humans. We call this an LLM-Modulo Framework (a name loosely inspired by SAT Modulo Theories Nieuwenhuis & Oliveras (2006)); see Figure 3. LLMs play a spectrum of roles in this architecture, from guessing candidate plans, to translating those plans into syntactic forms that are more accessible to external critics, to helping end users flesh out incomplete specifications, to helping expert users acquire domain models (that in turn drive model-based critics). All this leveraging of LLMs is done without ascribing to them any planning or verification abilities. The LLM ideas are vetted by external critics, thus ensuring that the plans generated in this architecture can have formal correctness guarantees where possible.
One question regarding this LLM-Modulo architecture from a planning perspective is whether it is more than a gratuitous attempt to shoe-horn LLMs to solve planning problems, when there are more formal combinatorial planning systems available in multiple communities Ghallab et al. (2004). Compared to a planner that is guaranteed to be correct in a narrow set of domains, LLMs may likely be good at generating plausible (but not guaranteed to be correct) plan heuristics/suggestions in many more scenarios. Thus, unlike the traditional planning architectures studied in AI Ghallab et al. (2004), which put a priori constraints on the expressiveness of the problems that can be posed to the planner (to wit, the different expressiveness levels of the PDDL specification McDermott et al. (1998)), the LLM-Modulo architecture puts no such restrictions. In this sense, it is more representative of real-world planning problems such as those in NASA mission planning, where the different critics–human and automated–are at best able to give “no objection” certificates, clearing it from their perspective. (Indeed, both deep space network planning and mars rover task planning are done via a collective human blackboard. Johnston et al. (2014) Bresina et al. (2004).) Note that this is starkly different from just sending an unvetted plan out to execution (as would be the case if we have LLMs operate in autonomous mode to guess plans). Generalizing planning and reasoning frameworks this way is consistent with the Doyle & Patil’s call to the Knowledge Representation community of yore Doyle & Patil (1991).
@@ 2.1 LLMs cannot generate executable plans in autonomous mode
They show that the results in the autonomous mode are pretty bleak. On average, only about 12% of the plans that the best LLM (GPT-4) generates are actually executable without errors and reach their goals. They show that the choice of the specific LLM (they have tested the family of GPT LLMs including GPT-4 OpenAI (2023), GPT-3.5 OpenAI (2022), InstructGPT-3 Ouyang et al. (2022) and GPT-3 Brown et al. (2020)). They also sho that fine-tuning does not seem to have a major effect on this dismal performance. They also show that the performance deteriorates further if the names of the actions and objects in the domain are obfuscated–a change that doesn’t in any way affect the performance of the standard AI planners. This latter further suggests that LLMs are more likely doing approximate retrieval of plans than planning.
@@ 2.2 LLMs cannot verify plans and thus cannot improve by self-critiquing
There still exists considerable optimism that even if LLMs can’t generate correct solutions in one go, their accuracy might improve in an iterative prompting regime, where LLMs will be able to “self-critique” their candidate solutions and refine them to the point of correctness Yao et al. (2023b, a); Shinn et al. (2023); Weng et al. (2023); Huang et al. (2022). This belief seems to rest largely on the assumption that verification of correctness should be easier than generation for many reasoning problems–a rather classical argument from computational complexity. There are grounds to be skeptical of this assumption as the complexity of the reasoning task should be irrelevant to LLM performance if what they are doing is approximate retrieval. In general, unless LLMs are trained not just on “correct data,” but also “corrections data,” there is no a priori reason to believe that their critiques would even be approximately relevant, let alone actually correct.
@@
Their results indicate that in direct mode, LLMs are, perhaps not surprisingly, pretty bad at solving graph coloring instances. More interestingly, they are no better at verifying solutions. In iterative modes, given the inability of LLMs to verify solutions, it should come as no surprise that their experiments also show that the strategy of LLMs self-critiquing their solutions does not improve over the baseline. They report that the perforance is in fact worse because the system can’t recognize a correct coloring and thus merrily passes over fortuitously correct colorings it has generated, ending up with a wrong one! Similar results have also been reported for planning problems in Valmeekam et al. (2023c).
One important corollary of the fact that LLMs cannot self-critique their plans is that they can’t also self-improve by generating synthetic data by generating plans themselves, critiquing the plans by themselves to improve them, and then using those to fine-tune themselves, as has been claimed in the literature Wang et al. (2022) [^3] Contrary to the claim of “self-improvement”, works like Wang et al. (2022) actually heavily depend on external knowledge (crafted seed examples) and critics (filtering step). Huang et al. (2023b); see Section 3.3.
@@ 2.3 Analyzing Claims to the Contrary in the Literature
Claims about Planning: To analyze the planning claims, we need to first understand that solving planning tasks requires (a) having the necessary planning domain knowledge–the actions and their preconditions, effects; the standard hierarchical recipes (e.g. task reduction schemas in HTN planning), past cases/plans, etc., and (b) being able to assemble this planning knowledge into an executable plan that takes care of any subgoal/resource interactions. The first can be called the knowledge acquisition and the second reasoning/planning part. Many of the papers claiming planning abilities of LLMs, on closer examination, wind up confusing general planning knowledge extracted from the LLMs for executable plans. When all we are looking for are abstract plans, such as “wedding plans,” with no intention of actually executing the said plans, it is easy to confuse them for complete executable plans. Indeed, our close examination of several works claiming planning capabilities for LLMs Kambhampati et al. (2023) suggests that they either work in domains/tasks where subgoal interactions can be safely ignored Yao et al. (2023b); Shinn et al. (2023) [^4] Although domains like AlfWorld Shridhar et al. (2021) do have sub-goal interactions for successful task completion, Yao et al. (2023b) and Shinn et al. (2023) ignore these interactions relying on the ergodic nature of the domain when prompting LLMs for generating plans. –either because they are just working on a single goal, or because the world is forgiving and ergodic; or delegate the interaction resolution (reasoning) to the humans in the loop (who, through repeated prompting, have to “correct” the plan). Sometimes, in common sense domains, or with enough fine-tuning, the “assembling” part may also be obviated by having seen a case that pretty much corresponds to the problem that needs to be solved. Not surprisingly, the work by Valmeekam et al. (2023c) shows that if the action interactions are removed by relaxing the world models, then the ability of LLMs to guess executable plans improves. Without these assumptions or mitigations, the plans that come out of LLMs may look reasonable to the lay user, and yet lead to execution time interactions and errors. [^5] These issues are illustrated in part by a recent news story Kugel & Hiltner (2023) about the proliferation of travel planning books, mostly auto-extracted from LLMs, and the ensuing disappointment of the unsuspecting end users who buy them mistaking them for usable plans!
@@
The fact that LLMs are often good at extracting planning knowledge can indeed be gainfully leveraged. As shown in recent works Guan et al. (2023), LLMs can be a rich source of approximate models of world/domain dynamics and user preferences, as long as the humans (and any specialized critics) in the loop verify and refine those models, and give them over to model-based solvers. This way of using LLMs has the advantage that the humans need only be present when the dynamics/preference model is being teased out and refined, and the actual planning after that can be left to sounder planning frameworks with correctness guarantees, such as LLM-Modulo framework we propose.
@@
Such an overall approach has striking similarities to knowledge-based AI systems of yore, with LLMs effectively replacing the “knowledge engineer” (see Figure 2). Given the rather quixotic and dogmatic shift of AI away from approaches that accept domain knowledge from human experts that some writers termed “Polanyi’s Revenge” (c.f. Kambhampati (2021)), this new trend of using LLMs as knowledge sources can be viewed as a form of avenging Polanyi’s revenge! Indeed, LLMs make it easy to get problem-specific knowledge as long as we are willing to relax the correctness requirements of that knowledge. In contrast to the old knowledge engineering approaches, LLMs offer this without making it look like we are inconveniencing any specific human (we are, instead, just leveraging everything humans told each other on the Web!). So the million dollar question for reasoning tasks is: “how would you do robust planning if you have some doddering know-it-all ready to give you any kind of knowledge?” The LLM-Modulo Framework is a principled method for tackling this challenge.
@@
In a related vein, there is the recent Tree of Thoughts (ToT) paper Yao et al. (2023a), which has been pitched as a way to convert LLMs into some type of systematic search with self-verification. A closer look at the work however shows that ToT simply iteratively back-prompts the LLM until it comes up with a solution that is acceptable to an external verifier. Specifically, ToT employs a problem-specific prompt priming method. The “tree” in ToT is essentially a way to generate diverse priming prompts (that the authors set up in a problem specific way). In other words, despite the use of terminology of problem-solving agents Russell & Norvig (2010)–search tree, expansion etc., there is really no deeper connection to search-based agents. [^7] Our preliminary experiments also show that at least in 24 puzzle, a simple iterative prompting, even without a systematic prompt diversification, is quite competitive with the ToT framework.
@@
In general planning problems, one way to provide an external verifier is to (a) write a domain model (e.g. in PDDL) and (b) feed it to an off-the-shelf model-based verifier like VAL (c.f. Howey et al. (2004)).
@@ 3.2 Backprompt (Meta) Controller
The processing steps taken in the controller can range from simple round-robin selection of prompts to generating a summarized prompt (with LLM help)111111 Such summarization is a reasonable strategy as the back prompts will not be treated as hard constraints by LLMs anyway. to employ a prompt diversification strategy to elicit the next candidate from a different part of the implicit search space (akin effectively to the strategy used in systems such as Tree of Thoughts prompting Yao et al. (2023a), as discussed in 2.3).
@@ 3.3 Fine Tuning & Synthetic Data
One important question is where this additional data for fine tuning comes from. A tempting idea is to have the LLM itself generate this additional data, and improve it by self-critiquing/verification, before fine-tuning itself on the data. This EM-like approach unfortunately will not work given that LLMs can’t verify their own solutions (see Section 2.2). In the past, this meant that the only reliable way to generate synthetic data is to use external plan generators–for example, use a classical planner like FF Hoffmann & Nebel (2001) to solve a host of blocks world instances and then use those solutions to fine-tune the LLM. The LLM-Modulo framework, in contrast, provides an alternative way of using an LLM-based framework to generate synthetic data that is guaranteed correct.
@@ 3.4 Specification Refinement & Model Acquisition (Semi-automated)
As mentioned earlier, we avoid having humans involved in iteratively prompting LLMs–as this can be an infeasibly time-consuming activity for them. Instead, we let automated verifiers, either model-based or LLM-supported, to manage the plan critiquing process. The framework does depend on humans for “once per domain” and “once per problem” interactions. In the former category, human domain experts can play a role in acquiring the domain model with the help of the LLM. Examples of such interaction include teasing out PDDL planning models from the LLMs with the help of human expert curation (top left in Figure 3). The idea here is that the traditional domain model acquisition task (e.g. sim (2001)) is significantly made easier by having the LLMs help with ideas regarding various pieces of the domain model (e.g., actions, their preconditions and effects) and letting humans sign off/critique the resulting model. Once the model is acquired this way, it can be used by correctness verifiers such as VAL Howey et al. (2004); Guan et al. (2023). Often the planning problems in real world situations are specified incompletely, leaving it to the human commonsense to refine the specification. This brings up a second role for humans–this time end users (bottom left in Figure 3–in collaboratively refining the specification with the help of LLMs (similar to the way done in Xie et al. (2023); Liu et al. (2023)).